OpenStudy (anonymous):

The sum of two consecutive integers is 141. Find the integers.

OpenStudy (anonymous):

Yeah, but I was hoping you could show me how to do the problem :)

OpenStudy (anonymous):

please..

OpenStudy (anonymous):

n + 1= 141 2n + 2 = 141 2n = 141 2n + 1 = 141

OpenStudy (shinalcantara):

i didn't see the 'consecutive' part. my bad -_-

OpenStudy (anonymous):

it's totally fine

OpenStudy (shinalcantara):

let 'x' be the 1st number 'x+1' be the 2nd number (since they are consecutive numbers) ----------------------- Having 141 as the sum you'll have it as: x +(x+1) = 141 ----------------------- Removing parenthesis x + x + 1 = 141 ----------------------- Transpose 1 to the right side of the equation x + x = 141 - 1 ----------------------- Combine like terms 2x = 140 ----------------------- Divide the whole equation with 2 x = 70

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

I need the equation :/ I suck at algebra

OpenStudy (shinalcantara):

trace the equation in the solution. The only difference is the variable used. i used 'x' while your choices is in 'n'

OpenStudy (anonymous):

fine, I totally suk at this and my brain isn't working this morning *gives gummy bear*

OpenStudy (shinalcantara):

Taken from above's solution: x + x + 1 = 141 combine like terms (add x and x) 2x + 1 = 141